OpenStudy (anonymous):

A parent died and left an estate to four children. One inherited 1/2 of the estate, the second inherited 1/8, and the third inherited 1/8. How much did the fourth inherit?

OpenStudy (anonymous):

thats soo sad lol

OpenStudy (anonymous):

yea it is lol...

OpenStudy (anonymous):

\[1-\frac{1}{2}-\frac{1}{8}-\frac{1}{8}\]

OpenStudy (anonymous):

you can do it in your head one step at a time one minus one half is one half one half minus one eighth it four eighths minus one eighth is three eighths three eighths minus one eighth is two eighths two eighths is one quarter

OpenStudy (anonymous):



OpenStudy (anonymous):

they got one quarter if the inheritance

OpenStudy (anonymous):

is it -0.75?

OpenStudy (anonymous):

no, you cannot inherit a negative amount of money

OpenStudy (anonymous):

look at the picture, i might be more clear

OpenStudy (anonymous):

Got it's 1/4